OpenStudy (anonymous):

You are told that P(A | B) = P(B | A). Which statement below must be true? A) A and B are equally likely B) P(A | B) = 0.5 C) A and B are independent D) A and B are dependent

ganeshie8 (ganeshie8):

@perl

OpenStudy (perl):

given P( A | B ) = P ( B | A ) this implies, by substituting P ( A & B ) / P(B) = P( B & A ) / P (A)

OpenStudy (perl):

Using the definition : P ( X | Y ) = P( X & Y ) / P ( Y )

OpenStudy (perl):

@AshleyLashae can you try to solve it now ? hint, cross multiply

OpenStudy (anonymous):

A

OpenStudy (perl):

and there is a commutativity rule you can apply as well: P( A & B) = P( B & A )